What is the (OH)- ion in the solution prepared by mixing 20 ml of .050 M HCl with 30ml of .10 M. Ba(OH)2?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
moles of OH- in BaOH2 initially is: = > 30 * 1/10 = > 3*2 = 6 mm
moles of H+ => 1 mm
i mm of OH- react with one mm of H+.
remaining mm of OH- => 6 – 1 = 5 mm.
ccn. = 5 x 10^-3 / 0.05 = 0.1 M